write a step by step instructions in english to calculate the fewest number of dollar bills needed to pay a bill of amount TOTAL.for examlpe,if TOTAL is $283,the bills would consist of two $100 bills,one $50 bill,one $20 bill,one $10 bill,one $5 bill,one $2 bill,and one $1 bill.(For this exercise assume that only $100,$50,$20,$10,$5,$2,$1 bills are available.)Do not write program in c++
1
Expert's answer
2012-03-01T09:35:39-0500
take ourstarting number, for example368.
1)divide 368 by100.368/100= 3.68,

we take only an integer - 3,

this means that the bills would consist of 3 - $100 bills,

thensubtract from 368 our 3 - $100 bills.368-300= 68.




2) divide 68 by50.68/50= 1.36,

we takeonly an integer - 1,

this means that the bills would consist of 1 - $50 bills,


thensubtract from 68 our 1 - $50 bills.68-50= 18.





3) divide 18 by20. 18/20= 0.9,

we have a number less then zero,
this means that the bills would not consist a $50 bills,

and we move on...



4) divide 18 by 10. 18/10 = 1.8,

we take only an integer - 1,

this means that the bills would consist of 1 - $10 bills,


then subtract from 18 our 1 - $10 bills. 18-10 = 8.





5) divide 8 by 5. 8/5 = 1.6,

we take only an integer - 1,

this means that the bills would consist of 1 - $5 bills,


then subtract from 8 our 1 - $5 bills. 8-5 = 3.




6) divide 3 by 2. 3/2 = 1.5,
we take only an integer - 1,
this means that the bills would consist of 1 - $2 bills,

then subtract from 3 our 1 - $2 bills. 3-2 = 1.



7) divide 1 by 1. 1/1 = 1,
we have an integer - 1,
this means that the bills would consist of 1 - $1 bills,
and this means that the algorithm is finished!